OpenStudy (anonymous):

Solve for x: |3x + 2| = 2

OpenStudy (anonymous):

dude, go to the website i gave you earlier. http://www.algebrahelp.com/calculators/equation/

OpenStudy (anonymous):

the answer is zero

OpenStudy (anonymous):

you mean no solution right

OpenStudy (anonymous):

3x+2=2 3x=0 x/3=0/3 x=0

OpenStudy (anonymous):

for positive charge we can write 3x+2=2 =>3x=2-2 =>3x=0 =>x=0 for negative charge we can write -(3x+2)=2 =>-3x-2=2 =>-3x=2+2 =>-3x=4 =>x=-4/3

OpenStudy (anonymous):

right. and well, he already gave you the next part